OpenStudy (marcelie):

find he position of a particle moving along a coordinate line that satisfies the given conditions a(t)= 2sint - cost s(0) = 6 v(0)=0

OpenStudy (xguardians):

a = acceleration (second derivative) s = position? I'm guessing? V = velocity (first derivative)

OpenStudy (marcelie):

hmm is that the same for anti derivatives ?

OpenStudy (anonymous):

\[a=\frac{ d^2s }{ dt^2 }\] \[v=\frac{ ds }{ dt }\] \[\frac{ d^2s }{ dt^2 }=2\sin t-\cos t\] integrate and find the value of constant from the given condition. Again integrate and find the value of constant from the given condition.